OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 + x - 20 A. x - 4 B. x + 2 C. x - 2 D. x + 4

OpenStudy (anonymous):

x-4

OpenStudy (anonymous):

x^2 + x - 20 = x^2 + 5x - 4x -20 = x(x+5) -4(x+5) = (x-4)(x+5) Can you get the answer now?

OpenStudy (anonymous):

thanks

OpenStudy (maheshmeghwal9):

\[\LARGE{\color{red}{x-4}.} \]